# Geometry-Informed Neural Operator Transformer

### v3

@@ Abstract
Machine-learning-based surrogate models offer significant computational efficiency and faster simulations compared to traditional numerical methods, especially for problems requiring repeated evaluations of partial differential equations. This work introduces the Geometry-Informed Neural Operator Transformer (GINOT), which integrates the transformer architecture with the neural operator framework to enable forward predictions for arbitrary geometries. GINOT encodes the surface points cloud of a geometry using a sampling and grouping mechanism combined with an attention mechanism, ensuring invariance to point order and padding while maintaining robustness to variations in point density. The geometry information is seamlessly integrated with query points in the solution decoder through the attention mechanism. The performance of GINOT is validated on multiple challenging datasets, showcasing its high accuracy and strong generalization capabilities for complex and arbitrary 2D and 3D geometries.
@@
GINOT uses a sampling, grouping, and attention mechanism to encode surface point clouds, ensuring invariance to point order and padding, and robustness to varying point densities.
@@ 1 Introduction
In this work, we introduce the Geometry-Informed Neural Operator Transformer (GINOT), which integrates the transformer architecture with the neural operator framework to enable forward predictions for arbitrary geometries without requiring SDF as geometry features. GINOT consists of two primary components: the geometry encoder and the solution decoder. Drawing inspiration from Shap-E [30] by OpenAI, we design the geometry encoder to process the boundary points cloud, representing the geometry, into the KEY and VALUE matrices for the attention mechanism. This choice facilitates the integration of information from query points in the solution decoder.
The geometry encoder employs sampling and grouping layers [31, 32] to extract local geometry information, which is then combined with global geometry information using cross-attention layers. This approach ensures invariance to point order and robustness to variations in point density. Since the number of points in the points cloud typically varies across different geometries, padding methods are used to ensure uniform input sizes within a batch. To prevent padding points from influencing the output of the geometry encoder, a masking mechanism is applied, which excludes padding points from selection in the sampling and grouping layers and ensures that they do not affect the attention scores in the cross-attention layer.
@@ 2 GINOT
Figure 1: Overview of the Geometry-Informed Neural Operator Transformer (GINOT) architecture. The boundary points cloud is initially processed through sampling and grouping layers to extract local geometric features. These local features are then fused with global geometric information via a cross-attention layer. This is followed by a series of self-attention layers and a final linear layer, producing the KEY and VALUE matrices for the cross-attention layer in the solution decoder. In the solution decoder, an multilayer perceptron (MLP) encodes the query points into the QUERY matrix for the cross-attention layer, which integrates the geometry information from the encoder. The output of the cross-attention layer is subsequently decoded into solution fields at the query points using another MLP.
@@ 2.1 Geometry encoder
To address these challenges, we developed a geometry encoder that processes the boundary points cloud of a geometry into the KEY and VALUE matrices for the attention mechanism. This enables seamless integration of geometry information with query points in the solution decoder. Boundary points clouds representing geometries exhibit three key characteristics: 1) invariance to the order of points, 2) non-uniform point density across different locations and geometries, and 3) varying numbers of points for different geometries. Drawing inspiration from Shap-E [30] by OpenAI, the geometry encoder, as shown in Fig. 1(a), is designed to handle these characteristics. It ensures invariance to point order and padding while maintaining robustness to variations in point density.
@@
The points cloud is first processed through a sampling layer and a grouping layer for local geometry features, originally introduced in PointNet [31] and PointNet++ [32], which are well-suited to handle the aforementioned properties. The sampling layer employs the iterative farthest point method [32] to select $N_{s}$ points from the points cloud. This method begins by randomly selecting a point, then iteratively selects the farthest point from the already selected points until $N_{s}$ points are chosen. These selected points are then passed to the grouping layer, which forms $N_{s}$ groups, each containing $N_{p}$ points.
@@
For each group, the $N_{p}$ points are selected as those within a ball of radius $r$ (a hyperparameter) centered at the corresponding sampled point. If fewer than $N_{p}$ points exist within the ball, the group is padded with the nearest points to the center. Conversely, if more than $N_{p}$ points are present, the nearest $N_{p}$ points are chosen. Consequently, the output of the grouping layer has dimensions $N_{s}\times N_{p}\times d$, where $d$ represents the dimensionality of the points cloud. The sampling and grouping layers ensure invariance to the order of points in the input points cloud. These layers output the selected points along with their indices in the original points cloud, which are subsequently used in downstream layers.
@@
The points cloud is also processed through a positional encoder layer utilizing the Nerf encoding method [33], followed by a linear layer. Using the indices from the sampling and grouping layers, the output is indexed to extract local geometry information, resulting in a tensor of shape $N_{s}\times N_{p}\times C$, where $C$ represents the output channel dimension of the positional encoder. This tensor is concatenated with the output of the grouping layer. The combined tensor is then passed through a series of 2D convolutional layers and a multi-layer perceptron (MLP), which aggregates the information within each group, reducing the dimension of $N_{p}$ to 1. As a result, the final output has a shape of $N_{s}\times C$.
@@
The local information obtained by the sampling and grouping layers is utilized as the QUERY ($Q$) in the cross-attention blocks, while the KEY ($K$) and VALUE ($V$) are derived from the output of the positional encoder layer, which captures global geometry information. The attention mechanism allows the local information in the QUERY ($Q$) to focus on the relevant global geometry features. The QUERY ($Q$) is invariant to the permutation of points due to the sampling and grouping mechanism, whereas the KEY ($K$) and VALUE ($V$) depend on the order of points as they are directly derived from the positional encoder layer with the points cloud as input. However, the output of the cross-attention block remains invariant to the order of points because of:
@@
The sampling and grouping layers, combined with the attention mechanism, ensure that the geometry encoder remains invariant to the permutation of points in the points cloud while maintaining robustness to variations in point density. Since the number of points in the points cloud varies across different geometries, padding is typically required to standardize input sizes within a batch for deep learning models. To prevent padding points from influencing the geometry encoder’s output, a masking mechanism is employed. This mechanism excludes padding points from being selected in the sampling and grouping layers. Additionally, padding points can interfere with attention scores in the cross-attention layer of the geometry encoder. To mitigate this, the corresponding entries in the attention score matrix $(Q\cdot K^{T})/\sqrt{d_{e}}$ are set to "$\infty$- ∞", ensuring that the attention layer’s output is unaffected by padding points. This approach guarantees that the geometry encoder is invariant to both point order and padding points, while also being robust to variations in points cloud density. The flexibility and robustness of the proposed geometry encoder are further illustrated in Section 3.5.
@@ 3.1 Overview
To evaluate the performance and accuracy of GINOT, six datasets are utilized: elasticity [22], Poisson equation with both structured and unstructured meshes, bracket lugs [24], micro-periodic unit cell (PUC) [34] in [9], and Jet Engine Bracket (JEB) [35]. Table 1 provides a detailed summary of these datasets, including their size, the number of query points, the number of points in the boundary points cloud, whether the dataset is parametric or fully arbitrary, and the type of solution considered.
@@ 3.2 Elasticity
Figure 3: Stress solutions of the testing sample for the elasticity problem. The first row shows the median case ($L_{2}=1.23\%$), while the second row represents the worst case ($L_{2}=2.93\%$) based on $L_{2}$ relative error. The first column depicts the input surface points cloud, the second column presents the ground truth from finite element analysis, the third column shows the prediction by GINOT, and the last column highlights the absolute error between the prediction and the ground truth.
@@ 3.3 Poisson equation
Figure 4: Testing sample solutions for the Poisson equation. The first row shows the median case ($L_{2}=0.41\%$) while the second row presents the worst case ($L_{2}=2.27\%$) for unstructured mesh data, based on $L_{2}$ relative error. The third row displays the median case ($L_{2}=0.38\%$) for structured mesh data. The first column illustrates the input surface points cloud, the second column shows the FEM ground truth, the third column provides GINOT predictions, and the last column highlights the absolute error between predictions and the ground truth. Structured and unstructured meshes are also depicted in the last column.
@@ 3.4 Bracket lug
We address the plasticity problem of bracket lugs using the dataset used in [24]. The bracket lugs, depicted in Fig. 5, are parameterized by three variables: length, thickness, and the radius of the circular hole. FEM simulations employ an elastic-plastic material model with linear isotropic hardening and quadratic hexahedral elements under small deformation assumptions. For training GINOT, the middle edge nodes of the quadratic hexahedral elements are removed. The surface mesh is extracted from the volume hexahedral elements, and its nodes are used as the boundary points cloud, which is then input to the geometry encoder of GINOT.
@@
Figure 5: Stress solutions for the testing samples of bracket lugs. The first row shows the median case ($L_{2}=0.41\%$), while the second row represents the worst case ($L_{2}=3.78\%$) based on $L_{2}$ relative error. The first column illustrates the input surface points cloud, the second column presents the ground truth from finite element analysis, the third column shows the GINOT prediction, and the last column highlights the absolute error between the prediction and the ground truth.
@@ 3.5 Micro-periodic unit cell
Figure 6: Mises stress and displacement solutions at the final strain step for the testing sample of the micro-periodic unit cell. The first row shows the median case ($L_{2}=10.37\%$), while the second row represents the 95th percentile case ($L_{2}=19.64\%$) based on $L_{2}$ relative error. The first column illustrates the input surface points cloud, the second column displays the true stress on the true deformed shape, the third column shows the predicted stress on the predicted deformed shape, and the last column highlights the absolute error of stress on the true deformed shape.
Figure 7: Demonstrating the versatility of the geometry encoder. The first row illustrates the input surface points cloud, with colors indicating the point order. The second row presents the corresponding predicted Mises stress on the predicted deformed shape by GINOT. (a) Original surface points cloud (PC) with 220 points, (b) shuffled original PC with the first 80% retained, (c) original PC padded with 199 points at coordinates (-1000, -1000) (not shown), and (d) shuffled version of the PC in (c). The $L_{2}$ relative errors for each case are displayed at the bottom.
@@
Periodic boundary conditions and compression are applied to the unit cell for FEM simulations to obtain the mechanical response field. The simulations use an elasto-plastic material model with large deformation within the environment of Abaqus. To approximate quasi-static compression while improving convergence for large deformation and elasto-plastic behavior, an implicit dynamic solver is employed with a virtual mass density of $\rho=10^{-8}$. The compression strain is applied smoothly over 51 steps from $\varepsilon=0$ to $\varepsilon=20\%$. The simulation utilizes a combination of mixed first-order quadrilateral elements and second-order triangular elements with full integration. To augment the dataset, the unit cell is randomly shifted multiple times. The corresponding stress and displacement solution fields are also shifted accordingly, enabling data augmentation without additional FEM simulations. The boundary points cloud of the geometry is formed by combining the points on the contours from the Marching Squares algorithm and the points on the boundary of the square domain, spaced at intervals of $\frac{1}{32}$ mm. This boundary points cloud represents the geometry features and is fed into the geometry encoder.
@@
The predicted solutions at the final strain step are shown in Fig. 6, comparing the ground truth with GINOT predictions. The first row highlights the median case with an $L_{2}$ relative error of 10.37%, while the second row shows the 95th percentile case with an $L_{2}$ relative error of 19.64%. The first column illustrates the input surface points cloud, the second column presents the true stress on the true deformed shape, the third column displays the predicted stress on the predicted deformed shape, and the last column shows the absolute error of stress on the true deformed shape. The mean $L_{2}$ relative error across the testing samples is 9.05%, with a standard deviation of 3.53%.
As detailed in Section 2.1, the geometry encoder of GINOT is highly versatile, being invariant to permutations and padding of the points cloud while maintaining robustness to variations in point density. To illustrate this versatility, we use the second-best test case. The points cloud is shuffled, reduced to 80% of its original size, padded with 199 points at coordinates (-1000, -1000), and then shuffled again. These modified points clouds are input into the trained GINOT, and the results are presented in Fig. 7. The $L_{2}$ relative errors for the original, reduced and shuffled, padded, and shuffled padded points clouds are 4.24%, 4.65%, 4.24%, and 4.57%, respectively. The close $L_{2}$ relative errors across these cases, with the padded points cloud matching the original, highlight the robustness of the geometry encoder. However, significantly reducing the points cloud density leads to a notable increase in the $L_{2}$ relative error, as critical geometry information is lost. In contrast, shuffling and padding have minimal impact on the $L_{2}$ relative error.
@@ 3.6 Jet engine bracket
Figure 11: Mises stress solutions for the testing samples of the JEB dataset. The first row shows the best case ($\text{NRMSE}=0.74\%$, $L_{2}=16.9\%$), the second row represents the median case ($\text{NRMSE}=1.62\%$, $L_{2}=31.2\%$), and the third row depicts the worst case ($\text{NRMSE}=3.54\%$, $L_{2}=56.8\%$) based on NRMSE. The first column illustrates the input surface points cloud, the second column presents the ground truth from finite element analysis, the third column shows the GINOT prediction, and the last column highlights the absolute error between the prediction and the ground truth.
The dataset includes CAD models, volume meshes in VTK format, and solution fields on fine quadratic tetrahedral elements. Instead of extracting surface nodes from the volume mesh, we extract the boundary points cloud directly from the CAD model vertices for each JEB sample using FreeCAD with PythonAPI, as illustrated in Fig. 9(a). To reduce computational costs, the volume mesh is downgraded from 10-node tetrahedrons to 4-node tetrahedrons by removing the middle nodes of each element’s edge. The nodes of these linear tetrahedrons are then used as query points in the solution decoder of GINOT, as shown in Fig. 9(a). Fig. 9(b) illustrates the number of boundary points in the cloud and the volume mesh nodes for the 2,138 JEB samples. The boundary points cloud contains between 2,253 and 10,388 points, while the volume mesh nodes range from 18,085 to 55,009.

### v4

@@ Abstract
Machine-learning-based surrogate models offer significant computational efficiency and faster simulations compared to traditional numerical methods, especially for problems requiring repeated evaluations of partial differential equations. This work introduces the Geometry-Informed Neural Operator Transformer (GINOT), which integrates the transformer architecture with the neural operator framework to enable forward predictions on arbitrary geometries. GINOT employs a sampling and grouping strategy together with an attention mechanism to encode surface point clouds that are unordered, exhibit non-uniform point densities, and contain varying numbers of points for different geometries. The geometry information is seamlessly integrated with query points in the solution decoder through the attention mechanism. The performance of GINOT is validated on multiple challenging datasets, showcasing its high accuracy and strong generalization capabilities for complex and arbitrary 2D and 3D geometries.
@@
GINOT encodes surface point clouds that are unordered, have non-uniform point density, and varying numbers of points.
@@ 1 Introduction
In this work, we introduce the Geometry-Informed Neural Operator Transformer (GINOT), which integrates the transformer architecture with the neural operator framework to enable forward predictions for arbitrary geometries without requiring SDF as geometry features. GINOT consists of two primary components: the geometry encoder and the solution decoder. Drawing inspiration from Shap-E [30] by OpenAI, we design the geometry encoder to process the boundary point cloud, representing the geometry, into the KEY and VALUE matrices for the attention mechanism. This choice facilitates the integration of information from query points in the solution decoder.
The geometry encoder utilizes sampling and grouping layers [31, 32] to extract local geometric features from the boundary point cloud. These local features are fused with global geometry information through cross-attention layers, enabling effective processing of unordered and non-uniformly distributed point clouds—unlike convolutional neural networks (CNNs), which operate on regular grids with uniform density. Because the number of points in the point cloud can vary between geometries, padding is applied to standardize input sizes within a batch. To ensure that padding points do not influence the geometry encoder’s output, a masking mechanism is employed: it prevents padding points from being selected in the sampling and grouping layers and excludes them from contributing to attention scores in the cross-attention layer.
@@ 2 GINOT
Figure 1: Overview of the Geometry-Informed Neural Operator Transformer (GINOT) architecture. The boundary point cloud is initially processed through sampling and grouping layers to extract local geometric features. These local features are then fused with global geometric information via a cross-attention layer. This is followed by a series of self-attention layers and a final linear layer, producing the KEY and VALUE matrices for the cross-attention layer in the solution decoder. In the solution decoder, an multilayer perceptron (MLP) encodes the query points into the QUERY matrix for the cross-attention layer, which integrates the geometry information from the encoder. The output of the cross-attention layer is subsequently decoded into solution fields at the query points using another MLP.
@@ 2.1 Geometry encoder
To address these challenges, we developed a geometry encoder that processes the boundary point cloud of a geometry into the KEY and VALUE matrices for the attention mechanism. This enables seamless integration of geometry information with query points in the solution decoder. In contrast to CNN inputs, boundary point cloud used to represent geometries have three key characteristics: (1) invariance to the order of points, (2) non-uniform point density across locations, and (3) a variable number of points across geometries. Drawing inspiration from Shap-E [30] by OpenAI, the geometry encoder, as shown in Fig. 1(a), is designed to handle these characteristics. It ensures invariance to point order and padding while maintaining robustness to variations in point density.
@@
The point cloud is first processed through a sampling layer and a grouping layer for local geometry features, originally introduced in PointNet [31] and PointNet++ [32], which are well-suited to handle the aforementioned properties. The sampling layer employs the iterative farthest point sampling (FPS) method [32] to select $N_{s}$ points from the point cloud. This method begins by randomly selecting a point, then iteratively selects the farthest point from the already selected points until $N_{s}$ points are chosen. Compared to random sampling, the FPS method provides better coverage of the entire point set for a given number of centroids. The selected points are then passed to the grouping layer, which forms $N_{s}$ groups, each containing $N_{p}$ points.
@@
For each group, the $N_{p}$ points are selected as those within a ball of radius $r$ (a hyperparameter) centered at the corresponding sampled point. If fewer than $N_{p}$ points exist within the ball, the group is padded with the nearest points to the center. Conversely, if more than $N_{p}$ points are present, the nearest $N_{p}$ points are chosen. Consequently, the output of the grouping layer has dimensions $N_{s}\times N_{p}\times d$, where $d$ represents the dimensionality of the point cloud. The sampling and grouping layers allow to extract local geometry features from nonuniform density point cloud [32]. These layers output the selected points along with their indices in the original point cloud, which are subsequently used in downstream layers.
@@
The point cloud is also processed through a positional encoder layer utilizing the Nerf encoding method [33], followed by a linear layer. Using the indices from the sampling and grouping layers, the output is indexed to extract local geometry information, resulting in a tensor of shape $N_{s}\times N_{p}\times C$, where $C$ represents the output channel dimension of the positional encoder. This tensor is concatenated with the output of the grouping layer. The combined tensor is then passed through a series of 2D convolutional layers and a multi-layer perceptron (MLP), which aggregates the information within each group, reducing the dimension of $N_{p}$ to 1. As a result, the final output has a shape of $N_{s}\times C$.
@@
The local information obtained by the sampling and grouping layers is utilized as the QUERY ($Q$) in the cross-attention blocks, while the KEY ($K$) and VALUE ($V$) are derived from the output of the positional encoder layer, which captures global geometry information. The attention mechanism allows the local information in the QUERY ($Q$) to focus on the relevant global geometry features. The QUERY ($Q$) can be invariant to the permutation of points order based on the sampling and grouping mechanism, whereas the KEY ($K$) and VALUE ($V$) depend on the order of points as they are directly derived from the positional encoder layer with the point cloud as input. However, the output of the cross-attention block remains invariant to the order of points because of:
@@
The sampling and grouping layers, combined with the attention mechanism, allow that the geometry encoder handle the point cloud with varying points density, while remaining robust to the permutation of points order. Since the number of points in the point cloud varies across different geometries, padding is typically required to standardize input sizes within a batch for deep learning models. To prevent padding points from influencing the geometry encoder’s output, a masking mechanism is employed. This mechanism excludes padding points from being selected in the sampling and grouping layers. Additionally, padding points can interfere with attention scores in the cross-attention layer of the geometry encoder. To mitigate this, the corresponding entries in the attention score matrix $(Q\cdot K^{T})/\sqrt{d_{e}}$ are set to "$\infty$- ∞", ensuring that the attention layer’s output is unaffected by padding points. The flexibility and robustness of the proposed geometry encoder are illustrated in experiments Section 3.
@@ 3.1 Overview
To evaluate the performance and accuracy of GINOT, six datasets are utilized: elasticity [22], Poisson equation with both structured and unstructured meshes, bracket lugs [24], micro-periodic unit cell (PUC) [34] in [9], and Jet Engine Bracket (JEB) [35]. Table 1 provides a detailed summary of these datasets, including their size, the number of query points, the number of points in the boundary point cloud, whether the dataset is parametric or fully arbitrary, and the type of solution considered. The first three datasets have a constant number of points in the boundary point cloud, while the last three datasets feature a varying number of points in the boundary point cloud. All datasets feature non-uniform point densities, as illustrated in the visualization results in the following subsections.
@@ 3.2 Elasticity
Figure 3: Stress solutions of the testing sample for the elasticity problem. The first row shows the median case ($L_{2}=1.23\%$), while the second row represents the worst case ($L_{2}=2.93\%$) based on $L_{2}$ relative error. The first column depicts the input surface point cloud, the second column presents the ground truth from finite element analysis, the third column shows the prediction by GINOT, and the last column highlights the absolute error between the prediction and the ground truth.
@@ 3.3 Poisson equation
Figure 4: Testing sample solutions for the Poisson equation. The first row shows the median case ($L_{2}=0.41\%$) while the second row presents the worst case ($L_{2}=2.27\%$) for unstructured mesh data, based on $L_{2}$ relative error. The third row displays the median case ($L_{2}=0.38\%$) for structured mesh data. The first column illustrates the input surface point cloud, the second column shows the FEM ground truth, the third column provides GINOT predictions, and the last column highlights the absolute error between predictions and the ground truth. Structured and unstructured meshes are also depicted in the last column.
@@ 3.4 Bracket lug
We address the plasticity problem of bracket lugs using the dataset used in [24]. The bracket lugs, depicted in Fig. 5, are parameterized by three variables: length, thickness, and the radius of the circular hole. FEM simulations employ an elastic-plastic material model with linear isotropic hardening and quadratic hexahedral elements under small deformation assumptions. For training GINOT, the middle edge nodes of the quadratic hexahedral elements are removed. The surface mesh is extracted from the volume hexahedral elements, and its nodes are used as the boundary point cloud, which is then input to the geometry encoder of GINOT.
@@
Figure 5: Stress solutions for the testing samples of bracket lugs. The first row shows the median case ($L_{2}=0.41\%$), while the second row represents the worst case ($L_{2}=3.78\%$) based on $L_{2}$ relative error. The first column illustrates the input surface point cloud, the second column presents the ground truth from finite element analysis, the third column shows the GINOT prediction, and the last column highlights the absolute error between the prediction and the ground truth.
@@ 3.5 Micro-periodic unit cell
Figure 6: Mises stress and displacement solutions at the final strain step for the testing sample of the micro-periodic unit cell. The first row shows the median case ($L_{2}=10.37\%$), while the second row represents the 95th percentile case ($L_{2}=19.64\%$) based on $L_{2}$ relative error. The first column illustrates the input surface point cloud, the second column displays the true stress on the true deformed shape, the third column shows the predicted stress on the predicted deformed shape, and the last column highlights the absolute error of stress on the true deformed shape.
@@
Figure 8: Demonstrating the versatility of the geometry encoder. The first row illustrates the input surface point cloud, with colors indicating the point order. The second row presents the corresponding predicted Mises stress on the predicted deformed shape by GINOT. (a) Original surface point cloud (PC) with 220 points, (b) shuffled original PC with the first 80% retained, (c) original PC padded with 199 points at coordinates (-1000, -1000) (not shown), and (d) shuffled version of the PC in (c). The $L_{2}$ relative errors for each case are displayed at the bottom.
@@
Periodic boundary conditions and compression are applied to the unit cell for FEM simulations to obtain the mechanical response field. The simulations use an elasto-plastic material model with large deformation within the environment of Abaqus. To approximate quasi-static compression while improving convergence for large deformation and elasto-plastic behavior, an implicit dynamic solver is employed with a virtual mass density of $\rho=10^{-8}$. The compression strain is applied smoothly over 51 steps from $\varepsilon=0$ to $\varepsilon=20\%$. The simulation utilizes a combination of mixed first-order quadrilateral elements and second-order triangular elements with full integration. To augment the dataset, the unit cell is randomly shifted multiple times. The corresponding stress and displacement solution fields are also shifted accordingly, enabling data augmentation without additional FEM simulations. The boundary point cloud of the geometry is formed by combining the points on the contours from the Marching Squares algorithm and the points on the boundary of the square domain, spaced at intervals of $\frac{1}{32}$ mm. This boundary point cloud represents the geometry features and is fed into the geometry encoder.
@@
The predicted solutions at the final strain step are shown in Fig. 6, comparing the ground truth with GINOT predictions. The first row highlights the median case with an $L_{2}$ relative error of 10.37%, while the second row shows the 95th percentile case with an $L_{2}$ relative error of 19.64%. The first column illustrates the input surface point cloud, the second column presents the true stress on the true deformed shape, the third column displays the predicted stress on the predicted deformed shape, and the last column shows the absolute error of stress on the true deformed shape. The mean $L_{2}$ relative error across the testing samples is 9.05%, with a standard deviation of 3.53%.
@@
As detailed in Section 2.1, the geometry encoder in GINOT is designed to handle 1) nonuniform point cloud densities, 2) permutations in point order, and 3) padding within the point cloud. The first property is evident throughout the examples in this work. To demonstrate the latter two properties, we analyze the second-best test case in this example. Notably, GINOT maintains stable inference performance even when the point cloud is made sparser. In this analysis, the point cloud is shuffled, reduced to 80% of its original size, padded with 199 points at $(-1000,-1000)$, and shuffled again. These modified point clouds are input to the trained GINOT, with results shown in Fig. 8. The $L_{2}$ relative errors for the original, reduced and shuffled, padded, and shuffled padded point clouds are 4.24%, 4.65%, 4.24%, and 4.57%, respectively. The similar $L_{2}$ errors across these cases demonstrate the geometry encoder’s robustness to point order and padding. However, a substantial reduction in point cloud density can increase the $L_{2}$ error due to loss of geometric information, while shuffling and padding have negligible impact.
@@ 3.6 Jet engine bracket
Figure 11: Mises stress solutions for the testing samples of the JEB dataset. The first row shows the best case ($\text{NRMSE}=0.74\%$, $L_{2}=16.9\%$), the second row represents the median case ($\text{NRMSE}=1.62\%$, $L_{2}=31.2\%$), and the third row depicts the worst case ($\text{NRMSE}=3.54\%$, $L_{2}=56.8\%$) based on NRMSE. The first column illustrates the input surface point cloud, the second column presents the ground truth from finite element analysis, the third column shows the GINOT prediction, and the last column highlights the absolute error between the prediction and the ground truth.
The dataset includes CAD models, volume meshes in VTK format, and solution fields on fine quadratic tetrahedral elements. Instead of extracting surface nodes from the volume mesh, we extract the boundary point cloud directly from the CAD model vertices for each JEB sample using FreeCAD with PythonAPI, as illustrated in Fig. 9(a). To reduce computational costs, the volume mesh is downgraded from 10-node tetrahedrons to 4-node tetrahedrons by removing the middle nodes of each element’s edge. The nodes of these linear tetrahedrons are then used as query points in the solution decoder of GINOT, as shown in Fig. 9(a). Fig. 9(b) illustrates the number of boundary points in the cloud and the volume mesh nodes for the 2,138 JEB samples. The boundary point cloud contains between 2,253 and 10,388 points, while the volume mesh nodes range from 18,085 to 55,009.